# ArmGS: Composite Gaussian Appearance Refinement for Modeling Dynamic Urban Environments

### v1

@@ Abstract
Figure 1: An illustration of our ArmGS for modeling urban scenes in autonomous driving. Our approach is capable of modeling fine-grained changes of background scenes and objects across frames and camera viewpoints.
1 Introduction
@@
Autonomous driving has made remarkable advancements in recent years, but how to effectively validate the safety and reliability of an autonomous driving system remains an open question. In real-world driving scenes, there are many corner cases, such as collisions, which cannot be easily collected, but are critical to the safety of an autonomous driving system. Simulation is one of the most useful ways to alleviate this problem. It enables the reconstruction and generation of various challenging driving scenes for downstream closed-loop evaluation.
@@
Traditionally, virtual-engine-based methods, such as CARLA [4], use classic graphics rendering systems for driving scene simulation, but the simulated scenes often have significant gaps from the real-world ones. On the other hand, data-driven simulation enables photorealistic driving scene modeling in a low-cost and efficient way [28, 29, 30, 27, 11]. With the great success of Neural Radiance Fields (NeRFs) for neural rendering, some researchers have adapted NeRFs for scene modeling in autonomous driving [29, 30, 27, 19]. Although NeRF-based methods have shown great potential for reconstructing realistic driving scenes, they mostly suffer from long training time and fail to achieve real-time rendering. More recently, some researchers have resorted to 3D Gaussian Splatting (3DGS) [12] to learn explicit 3D scene representations for driving scene modeling [28, 11, 21, 2, 35]. Due to the efficiency of explicit representations and differentiable tile rasterizers, 3DGS-based methods have shown promising results for high-fidelity reconstruction and real-time rendering. However, existing approaches for driving scene modeling mostly neglect to model fine-grained driving scene variations between frames and camera viewpoints [^1] Throughout this work, camera viewpoints refer to different locations of the same camera or different cameras. In autonomous driving simulation, scene images are captured in the wild along the movement of the ego vehicle. This unavoidably brings significant appearance changes of scenes and objects across frames and camera viewpoints due to lighting and camera exposure. Thus, failing to model these variations can lead to missing fine-grained details and suboptimal simulation results.
@@
In this work, we propose to exploit composite 3D Gaussian Splatting with Appearance RefineMent (ArmGS) from multiple granularities for dynamic urban environments modeling in autonomous driving. Our approach constructs a composite driving scene model based on 3DGS representations and employs a multi-level appearance modeling scheme to optimize a set of transformation parameters for composite Gaussian refinement. Specifically, we represent a driving scene with composite 3DGS and refine them at local Gaussian level, global image level and dynamic actor level, respectively. For local Gaussian level refinement, we extract latent Gaussian appearance representation to learn Gaussian-wise transformation parameters for Gaussian appearance refinement. This helps to learn local fine-grained appearance variations in the 3D Gaussian space. For global image level refinement, we extract latent image appearance representation to learn global image-wise transformation parameters for image color refinement. This helps to learn global image-wise appearance changes across frames and camera viewpoints. Moreover, to model dynamic variations of moving road actors (such as vehicles), we employ a light-weight spatial-temporal deformation model to transform dynamic actor Gaussians across time. This multi-level appearance modeling scheme allows for refining the composite driving scene Gaussians at multiple granularities. Consequently, scene and road actors appearance variations are collectively modeled to reconstruct richer fine-grained cues of driving scenes. Our experiments on multiple challenging autonomous driving datasets, namely, Waymo [23], KITTI [8], NOTR [29] and VKITTI2 [1], for driving scene modeling and novel view synthesis show that our approach not only reconstructs global scene appearance variations between frames and camera viewpoints, but also models local fine-grained changes of background and dynamic objects. An illustration is shown in Fig. 1. In summary, our contributions are:
  • Our method is the first method that proposes to explicitly embed multi-granularity appearances to model fine-grained changes for dynamic urban environments modeling. This fills a gap left by both existing NeRF and Gaussian splatting methods.
  • Our method uses a set of transformation parameters for composite Gaussian refinement, which is simple and effective and will not break the differentiability of splatting.
  • Our experiments on Waymo, KITTI, NOTR and VKITTI2 show the superiority of our method over the state-of-the-art methods.
@@
Figure 2: An overview of our ArmGS approach. Our approach refines composite driving scene Gaussians with appearance modeling at multiple granularities, ranging from local Gaussians level to global images level and dynamic actors level. The modules for local level refinement, global level refinement and actor level refinement are indicated in yellow, green and blue, respectively.
2 Related Work
3D Gaussian Splatting.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) [12] is a recently proposed real-time rendering approach. Unlike NeRFs that employ a continuous radiance field to implicitly represent 3D scenes, 3DGS represents a 3D scene with a collection of explicit 3D Gaussian primitives and utilizes splatting-based rasterization [36] to project 3D Gaussians to 2D image space. Since the success of [12], researchers have made many efforts to improve 3DGS from different aspects [31, 5, 32, 3, 16], such as anti-aliasing [31], compression [5], surface reconstruction [16], etc. Among them, Dahmani et al. [3] propose to model image appearance changes with image embeddings for novel-view synthesis from unconstrained image collections. Our approach significantly differs from [3] in that we devise a novel multi-level appearance modeling scheme to optimize a set of transformation parameters for composite Gaussian refinement, rather than modeling transient objects and landmarks from unconstrained images. In addition, Wu et al. [26] propose to explore deformation fields with HexPlane representations for real-time dynamic 4D scene rendering, which is subsequently modified by [11] for dynamic urban scene reconstruction. Although we also employ deformation fields for dynamic actor level refinement, our method is different in that we construct a light-weight network structure and encode position, timestamp and spherical harmonics to learn actor variations, rather than constructing HexPlane representation [26, 11].
Autonomous Driving Scene Modeling.
@@
Autonomous driving simulation [4, 30, 28, 29, 15, 14] is critical for real-world closed-loop evaluation. For dynamic driving scene modeling, there are primarily two types of methods, namely, virtual-engine-based methods [4, 1] and data-driven methods [28, 29, 30, 11]. While virtual-engine-based methods typically lack scalability and suffer from significant domain gaps, data-driven methods are becoming prevailing in recent years due to their photo-realism, scalability and efficiency. For data-driven simulation, NeRF-based methods [7, 29, 19, 9] have shown compelling performance, but require point dense sampling along each ray to model the interaction between light and scene elements. This leads to expensive computational costs and low rendering speed. On the other hand, 3DGS-based methods [28, 6, 11, 21, 35] make full use of the advantages of 3DGS and achieve real-time rendering for modeling driving scenes. Yan et al. [28] proposes to model dynamic urban scenes with road asset pose optimization and 4D spherical harmonics appearance model, but neglects to model fine-grained driving scene variations between frames and camera viewpoints. Huang et al. [11] extends dynamic 4DGS [26] for driving scene reconstruction in a self-supervised manner, but the rendering results are suboptimal with some artifacts and do not explicitly model fine-grained changes of driving scenes. Our work also builds on 3DGS, but different from existing works, our approach refines composite driving scene Gaussians with appearance modeling at multiple granularities, ranging from local Gaussians to global images and dynamic actors. Our approach is capable of modeling global scene appearance variations as well as local fine-grained changes of background and dynamic objects in autonomous driving scenes.
3 Methodology
3.1 Preliminaries
3D Gaussian Splatting.
@@ 3.1 Preliminaries
Composite Driving Scene Gaussians.
@@
3.2 Composite Driving Gaussian with Multi-Level Appearance Refinement
@@
Local Composite Gaussian Level Refinement.
@@
$c^{\prime}=\alpha^{l}c+\beta^{l},\quad\text{where }(\alpha^{l},\beta^{l})=\mathcal{D}^{l}(f^{l}).$ (4)
$c^{\prime}$ is then used in Eq.(2) for color computation. In this way, the contribution of each Gaussian to the rendered image is modulated across frames, so fine-grained scene variations are effectively encoded into local Gaussian-wise appearance variations across frames. During rendering, to obtain a low-dimensional embedding $\epsilon$ for a novel view, we consider that the appearance variation of a novel view should be consistent with its adjacent frame. Thus, for each novel view, we calculate the index of its nearest neighbor training frames based on camera index and timestamp to get its corresponding $\epsilon$, and then perform Eqs.(3) and (4) for the transformation.
Global Image Level Refinement.
Although the local Gaussian-wise appearance refinement encourages each Gaussian to learn local fine-grained variations, it does not encourage global consistent variations across camera viewpoints. For example, image-wise appearance variation can be consistently changed due to direct sunlight or camera exposure, which requires the holistic scene appearance refinement. Thus, to compensate for the global image-wise variation, we further refine driving scene appearance at the global image level. Specifically, we encode camera location and viewing direction as a camera viewpoint code $\phi$ and merge it with the embedding $\epsilon$ as the latent image appearance representation $f^{g}$, which is defined as:
@@
$C^{\prime}=\alpha^{g}C+\beta^{g},\quad\text{where }(\alpha^{g},\beta^{g})=\mathcal{D}^{g}(f^{g}).$ (6)
With the collaboration of local Gaussian-wise appearance modeling and global image-wise appearance modeling, the composite driving scene Gaussians are modulated at different granularities.
Dynamic Actor Level Refinement.
In dynamic driving scene modeling, dynamic road actors usually have more complex motion and appearance variations than background and static objects. For example, a moving vehicle needs to be placed at different locations along its trajectory, while its appearance is affected not only by the scene but also by its actions, such as using brake signals. To deal with this problem, we also refine composite driving Gaussians at the dynamic actor level. Specifically, we employ a low-dimensional learnable embedding for each class of dynamic actors and construct a light-weight class encoder to learn the weights of the dynamic actor Hash grid, inspired by [10, 30]. This helps to encode class information into Hash grid for position encoding. Alternatively, to reduce training parameters, we observe that the simple yet effective sinusoidal encoding [17] also works well. Then, to model the complex appearance variations specific to each dynamic actor, we construct a light-weight spatial-temporal encoder and a light-weight deformation head to transform position and spherical harmonics attributes of dynamic actor Gaussians across time. Instead of using the HexPlane [26] to learn the spatial-temporal structure, we encode position $\mu$ with the encoding function $\mathcal{F}^{a}$ (class-wise Hash grid or sinusoidal encoding), timestamp $t$ with a sinusoidal function $\mathcal{F}^{f}$ and merge them with the spherical harmonics $h$ to learn the spatial-temporal actor representation $f^{a}$ of the dynamic actor. This is defined as:
@@
$\{\mu{{}^{\prime}},h{{}^{\prime}}\}{=}\{\mu{+}\Delta{\mu},h{+}\Delta{h}\},~{}\text{where }\{\Delta{\mu},\Delta{h}\}{=}\mathcal{D}^{h}(f^{a}).$ (8)
Note that, although Gaussian deformation is inspired by [26], our design differs from existing works [26, 11, 6] in that we construct light-weight encoders and deformation heads and encodes position, timestamp and spherical harmonics to learn actor variations, rather than constructing HexPlane representation [26, 11] or using neural fields for color computation [6]. Moreover, the collaboration of the local level refinement, the global level refinement and the dynamic actor level refinement brings superior performance for dynamic urban environments modeling in autonomous driving.
3.3 Composite Driving Gaussian Optimization
Training Objective.
@@ 3.3 Composite Driving Gaussian Optimization
Actor Pose Optimization.
@@
4 Experiment
4.1 Datasets and Evaluation Protocol
Datasets.
We conduct experiments on four challenging autonomous driving datasets, including the Waymo [23, 28], KITTI [8], NOTR [29] and Virtual KITTI 2 (VKITTI2) [1]. On Waymo, following [28], we select eight challenging sequences with dynamic actors and complex background and use the tracked boxes provided by [28] for experiments. We select every fourth frame as the testing novel view and use the remaining frames as the training views. We set the image resolution to $1066{\times}1600$ and evaluate both the novel view synthesis setting with the testing views and the image reconstruction setting with the training views. On KITTI, following [24, 7], we select three sequences and use the official tracklets for evaluation in the novel view synthesis setting. We select every second frame as the testing novel view while the others are used for training. The image resolution is set to $375{\times}1242$. On NOTR, following [29], we employ two splits, namely, the static-32 and dynamic-32 splits, which consist of 64 multi-camera driving scene sequences with various challenges, e.g., cross-camera background variations, various dynamic actors, etc. The image resolution is set to $640{\times}960$. On VKITTI2, following [34], we select two sequences for experiments and use the $50\%$ dropout rate evaluation protocol. The image resolution is set to $375{\times}1242$.
Evaluation Metrics.
Following [28, 29, 11, 34], we use PSNR, SSIM [25] and LPIPS [33] as the evaluation metrics.
Method Waymo Novel View Waymo Reconstruction KITTI
@@ 4.1 Datasets and Evaluation Protocol
StreetGaussians [28] 34.6 0.938 0.079 36.3 0.948 0.074 27.8 0.889 0.069
ArmGS (Ours) 35.7 0.944 0.074 38.1 0.957 0.064 30.3 0.931 0.036
Table 1: Quantitative comparison with the state-of-the-art methods on the Waymo and KITTI datasets.
Figure 3: Qualitative comparison with the state-of-the-arts on the Waymo dataset. From the first column to the fourth column, we show results for GT, S3Gaussian, StreetGaussians and Ours, respectively. We highlight some fine-grained details, e.g., traffic lights, vehicles and trees.
Figure 4: Qualitative comparison with the state-of-the-art methods on the KITTI dataset. From the first column to the 3th column, we show results for GT, StreetGaussians and Ours, respectively. We highlight some fine-grained details, e.g., vehicles and lane lines.
Figure 5: Qualitative comparison on the NOTR dataset. We highlight some fine-grained details, e.g., vehicles and pedestrians.

### v2

@@ Abstract
I INTRODUCTION
@@
Autonomous driving has made remarkable advancements in recent years, but how to effectively validate the safety and reliability of an autonomous driving system remains an open question. In real-world scenarios, there are many corner cases which cannot be easily collected but are critical to the safety of autonomous driving. Simulation is one of the most useful ways to alleviate this problem. It enables the reconstruction and generation of various challenging driving scenes for downstream closed-loop evaluation.
Figure 1: An illustration of our ArmGS for modeling urban scenes. Our approach is capable of modeling fine-grained changes of background scenes and objects across frames and camera viewpoints.
@@
Traditionally, virtual-engine-based methods, such as CARLA [5], use classic graphics rendering systems for driving scene simulation, but the simulated scenes often have significant gaps from the real-world ones. On the other hand, data-driven simulation enables photorealistic driving scene modeling in a low-cost and efficient way [31, 32, 3, 13, 33, 30]. With the great success of Neural Radiance Fields (NeRFs) for neural rendering, some researchers have adapted NeRFs for scene modeling in autonomous driving [32, 33, 30, 21]. Although NeRF-based methods have shown great potential for reconstructing realistic driving scenes, they mostly suffer from long training time and fail to achieve real-time rendering. More recently, some researchers have resorted to 3D Gaussian Splatting (3DGS) [14] to learn explicit 3D scene representations for driving scene modeling [31, 13, 3, 18, 2, 38]. Due to the efficiency of explicit representations and differentiable tile rasterizers, 3DGS-based methods have shown promising results for high-fidelity reconstruction and real-time rendering. However, existing approaches for driving scene modeling mostly neglect to model fine-grained variations across frames and camera viewpoints. In autonomous driving simulation, scene images are captured in the wild along the movement of the ego vehicle. This unavoidably brings significant appearance changes of scenes and objects across frames and camera viewpoints due to lighting and camera exposure. Thus, failing to model these variations can lead to the loss of fine-grained details and suboptimal results.
@@
In this work, we propose to exploit composite 3D Gaussian Splatting with Appearance RefineMent (ArmGS) from multiple granularities for dynamic urban environments modeling in autonomous driving. Our approach constructs a composite driving scene model based on 3DGS representations and employs a multi-level appearance modeling scheme to optimize a set of transformation parameters for composite Gaussian refinement. Specifically, we represent a driving scene with composite 3DGS and refine 3DGS at local Gaussian level, global image level and dynamic actor level, respectively. For local Gaussian level refinement, we extract latent Gaussian appearance representation to learn Gaussian-wise transformation parameters for Gaussian appearance refinement. This helps to learn local fine-grained appearance variations in the 3D Gaussian space. For global image level refinement, we extract latent image appearance representation to learn global image-wise transformation parameters for image color refinement. This helps to learn global image-wise appearance changes across frames and camera viewpoints. Moreover, to model dynamic variations of moving road actors (such as vehicles), we employ a light-weight spatial-temporal deformation model to transform dynamic actor Gaussians across time. This multi-level appearance modeling scheme allows for refining the composite driving scene Gaussians at multiple granularities. Consequently, scene and road actors appearance variations are collectively modeled to reconstruct richer fine-grained cues of driving scenes. Our experiments on multiple challenging autonomous driving datasets, namely, Waymo [25], KITTI [10], NOTR [32] and VKITTI2 [1], for driving scene modeling and novel view synthesis show that our approach not only reconstructs global scene appearance variations between frames and camera viewpoints, but also models local fine-grained changes of background and dynamic objects. An illustration is shown in Fig. 1. In summary, our contributions are: (I) Our method is the first method that proposes to explicitly embed multi-granularity appearances to model fine-grained changes for dynamic urban environments modeling, which fills a gap left by existing methods; (II) Our method employs a set of transformation parameters for composite Gaussian refinement, which is simple, effective, and preserves the differentiability of the splatting process; (III) Our thorough experiments on Waymo, KITTI, NOTR and VKITTI2 show the superiority of our method over the state-of-the-art methods.
@@
Figure 2: An overview of our approach. Our approach refines composite driving scene Gaussians with appearance modeling at multiple granularities, ranging from local Gaussians level to global images level and dynamic actors level. The modules for local level refinement, global level refinement and actor level refinement are indicated in yellow, green and blue, respectively.
II RELATED WORK
II-A 3D Gaussian Splatting
@@ II RELATED WORK
3D Gaussian Splatting (3DGS) [14] is a recently proposed real-time rendering approach. Unlike NeRFs that employ a continuous radiance field to implicitly represent 3D scenes, 3DGS represents a 3D scene with a collection of explicit 3D Gaussian primitives and utilizes splatting-based rasterization [39] to project 3D Gaussians to 2D image space. Since the success of [14], researchers have made many efforts to improve 3DGS from different aspects [7, 34, 35, 4, 22], such as anti-aliasing [34], compression [6], surface reconstruction [17], etc. Among them, SWAG [4] proposes to model image appearance changes with image embeddings for novel-view synthesis from unconstrained image collections. Our approach significantly differs from [4] in that we devise a novel multi-level appearance modeling scheme to optimize a set of transformation parameters for composite Gaussian refinement, rather than modeling transient objects and landmarks from unconstrained images. In addition, 4DGS [29] proposes to explore deformation fields with HexPlane representations for real-time dynamic 4D scene rendering, which is subsequently modified by [13] for dynamic urban scene reconstruction. Although we also employ deformation fields for dynamic actor level refinement, our method is different in that we construct a light-weight network structure and encode position, timestamp and spherical harmonics to learn actor variations, rather than constructing HexPlane representation [29, 13].
II-B Autonomous Driving Scene Modeling
@@
Autonomous driving simulation [5, 33, 31, 32, 15, 16] is critical for real-world closed-loop evaluation. For dynamic driving scene modeling, there are primarily two types of methods, namely, virtual-engine-based methods [5, 1] and data-driven methods [31, 32, 33, 18]. While virtual-engine-based methods typically lack scalability and suffer from significant domain gaps, data-driven methods are becoming prevailing in recent years due to their photo-realism, scalability and efficiency. For data-driven simulation, NeRF-based methods [9, 32, 21, 11] have shown compelling performance, but require point dense sampling along each ray to model the interaction between light and scene elements. This leads to expensive computational costs and low rendering speed. On the other hand, 3DGS-based methods [31, 13, 3, 8, 24] make full use of the advantages of 3DGS and achieve real-time rendering for modeling driving scenes. StreetGS [31] proposes to model dynamic urban scenes with road asset pose optimization and 4D spherical harmonics appearance model, but neglects to model fine-grained driving scene variations between frames and camera viewpoints. S3Gaussian[13] extends dynamic 4DGS [29] for driving scene reconstruction in a self-supervised manner, but the rendering results are suboptimal with some artifacts and do not explicitly model fine-grained changes of driving scenes. Our work also builds on 3DGS, but different from existing works, our approach refines composite driving scene Gaussians with appearance modeling at multiple granularities, ranging from local Gaussians to global images and dynamic actors. Our approach is capable of modeling global scene appearance variations, local fine-grained changes of background and dynamic objects in autonomous driving scenes.
III METHODOLOGY
III-A Preliminaries
3D Gaussian Splatting
@@ III-A Preliminaries
Composite Driving Scene Gaussians
@@
III-B The Proposed Approach
@@
Local Composite Gaussian Level Refinement
@@
$c^{\prime}=\alpha^{l}c+\beta^{l},$ (4)
where $(\alpha^{l},\beta^{l})=\mathcal{D}^{l}(f^{l})$. $c^{\prime}$ is then used in Eq.(2) for color computation. In this way, the contribution of each Gaussian to the rendered image is modulated across frames, so fine-grained scene variations are effectively encoded into local Gaussian-wise appearance variations across frames. During rendering, to obtain a low-dimensional embedding $\epsilon$ for a novel view, we consider that the appearance variation of a novel view should be consistent with its adjacent frame. Thus, for each novel view, we calculate the index of its nearest neighbor training frames based on camera index and timestamp to get its corresponding $\epsilon$, and then use Eqs.(3) and (4) for the transformation.
Global Image Level Refinement
Although the local Gaussian-wise appearance refinement encourages each Gaussian to learn local fine-grained variations, it does not encourage global consistent variations across camera viewpoints. For example, image-wise appearance variation can be consistently changed due to direct sunlight or camera exposure, which cannot be resolved by image preprocessing and requires the holistic scene appearance refinement. Thus, to compensate for the global image-wise variation, we further refine driving scene appearance at the global image level. Specifically, we encode camera location and viewing direction as a camera viewpoint code $\phi$ and merge it with the embedding $\epsilon$ as the latent image appearance representation $f^{g}$, which is defined as:
@@
$C^{\prime}=\alpha^{g}C+\beta^{g},$ (6)
where $(\alpha^{g},\beta^{g})=\mathcal{D}^{g}(f^{g})$. With the collaboration of local Gaussian-wise appearance modeling and global image-wise appearance modeling, the composite driving scene Gaussians are modulated at different granularities.
Dynamic Actor Level Refinement
In dynamic driving scene modeling, dynamic road actors usually have more complex motion and appearance variations than the background and static objects. For example, a moving vehicle needs to be placed at different locations along its trajectory, while its appearance is affected not only by the scene but also by its actions, such as using brake signals. To deal with this problem, we also refine composite driving Gaussians at the dynamic actor level. Specifically, we employ a low-dimensional learnable embedding for each class of dynamic actors and construct a light-weight class encoder to learn the weights of the dynamic actor Hash grid, inspired by [12, 33]. This helps to encode class information into the Hash grid for position encoding. Alternatively, to reduce training parameters, we observe that the simple yet effective sinusoidal encoding [19] also works well. Then, to model the complex appearance variations specific to each dynamic actor, we construct a light-weight spatial-temporal encoder and a light-weight deformation head to transform position and spherical harmonics attributes of dynamic actor Gaussians across time. Instead of using the HexPlane [29] to learn the spatial-temporal structure, we encode position $\mu$ with the encoding function $\mathcal{F}^{a}$ (class-wise Hash grid or sinusoidal encoding), timestamp $t$ with a sinusoidal function $\mathcal{F}^{f}$ and merge them with the spherical harmonics $h$ to learn the spatial-temporal actor representation $f^{a}$ of the dynamic actor. This is defined as:
@@
$\{\mu{{}^{\prime}},h{{}^{\prime}}\}{=}\{\mu{+}\Delta{\mu},h{+}\Delta{h}\},$ (8)
where $\{\Delta{\mu},\Delta{h}\}{=}\mathcal{D}^{h}(f^{a})$. Note that, although Gaussian deformation is inspired by [29], our design differs from existing works [29, 13, 8] in that we construct light-weight encoders and deformation heads and encodes position, timestamp and spherical harmonics to learn actor variations, rather than constructing HexPlane representation [29, 13] or using neural fields for color computation [8]. Moreover, the collaboration of the local level refinement, the global level refinement and the dynamic actor level refinement brings superior performance for dynamic urban environments modeling in autonomous driving.
III-C Composite Driving Gaussian Optimization
Training Objective
@@ III-C Composite Driving Gaussian Optimization
Actor Pose Optimization
@@
TABLE I: Quantitative comparison with the state-of-the-art methods on Waymo and KITTI.
Method Waymo Reconstruction Waymo Novel View KITTI
@@
StreetGS [31] 36.3 0.948 0.074 34.6 0.938 0.079 27.8 0.889 0.069
OmniRe [3] 35.9 0.953 0.108 31.9 0.899 0.126 - - -
SplatFlow [26] - - - - - - 27.9 0.927 0.093
ArmGS (Ours) 38.1 0.957 0.064 35.7 0.944 0.074 30.3 0.931 0.036
Figure 3: Qualitative comparison with the state-of-the-arts on Waymo. From the first to the fourth row, we show results of scene modeling under foggy, sunny, cloudy, and rainy conditions. We highlight some fine-grained details, e.g., traffic lights, vehicles and trees.
IV EXPERIMENT
IV-A Datasets and Evaluation Protocol
Datasets
We conduct experiments on four challenging autonomous driving datasets, including the Waymo [25, 31], KITTI [10], NOTR [32] and Virtual KITTI 2 (VKITTI2) [1]. On Waymo, following [31], we select eight challenging sequences with dynamic actors and complex background conditions and use the tracked boxes provided by [31] for experiments. We select every fourth frame as the testing novel view and use the remaining frames as the training views. We set the image resolution to $1066{\times}1600$ and evaluate both the novel view synthesis setting with the testing views and the image reconstruction setting with the training views. On KITTI, following [27, 9], we select three sequences and use the official tracklets for evaluation in the novel view synthesis setting. We select every second frame as the testing novel view while the others are used for training. The image resolution is set to $375{\times}1242$. On NOTR, following [32], we employ two splits, namely, the static-32 and dynamic-32 splits, which consist of 64 multi-camera driving scene sequences with various challenges, e.g., cross-camera background variations, various dynamic actors, etc.. The image resolution is set to $640{\times}960$. On VKITTI2, following [37], we select two sequences for experiments and use the $50\%$ dropout rate evaluation protocol. The image resolution is set to $375{\times}1242$.
Evaluation Metrics
Following [31, 32, 13], we use PSNR, SSIM [28] and LPIPS [36] as the evaluation metrics.